A trapezium whose parallel sides are 42m and 30m and the other sides are 18m and 18m . Find its area.

Grade:8

1 Answers

Vikas TU
6 years ago
 In trapezium ABCD, draw CE paralle to DA. 
Presently CE = 18 m 
Since, DC = 30 m in this way, AE = 30 m 
Likewise, EB = AB - AE = 42 - 30 = 12 m 
Presently, in ∆ EBC 
S = (18 + 18 + 12)/2 
= 48/2 
= 24 
= √(24 × 6 x 6 x 12) 
= 101.8 m² 
Draw CP perpendicular to EB. 
Range of ∆EBC = 1/2 × EB × CP 
101.8 = 1/2 × 12 × CP 
101.8/6 = CP 
⇒ CP = 17 m Therefore, h = 17 m 
Presently, range of trapezium = 1/2(p₁ + p₂) × h 
= 1/2 × 72 × 17 
= 612 m²